# Visual Writing: Writing by Manipulating Visual Representations of Stories

### v1

@@ 5.4. Results
Participants completed the task and the free-form exploration without major difficulties. Below, we summarize the quantitative findings using descriptive statistics. Following previous recommendations for study designs similar to ours (Zhu and Kolassa, 2018; Masson et al., 2023), confidence intervals are calculated using the studentized bootstrapping method using 10,000 repetitions.
Success rate
On a 5-point semantic differential scale from 1-unsuccessful to 5-successful, participants rated their success M=4.25 (95% CI: [3.8, 4.6]) at accomplishing the tasks. Specifically, they were most successful at the task about adding and removing entities (M=4.5, 95% CI: [3.8, 5]), followed by changing the locations of entities (M=4.25, 95% CI: [3.4, 4.8]), and reordering events (M=4, 95% CI: [2.8, 4.7]).
Creativity Support Index (CSI)
Participants generated an average CSI score of 71.5 (SD=21.12) for the system and the tasks. Table 1 shows the breakdown of the CSI results. Enjoyment was both the most highly rated factor and also the factor considered most important by participants. Immersion was ranked second most important yet received the lowest score. Possible reasons for these ratings are discussed in Section 6.
@@
Workload (Raw NASA-TLX)
Participants reported high performance (M=7.49 SD=1.95) for low mental demand (M=4.11, SD=2.38), effort (M=3.39, SD=2.17), frustration (M=2.5, SD=1.2), temporal demand (M=1.18, SD=1.04), and physical demand (M=1.18, SD=1.04).
Questionnaire
Overall, answers to the 5-point scale statements from 1-strongly disagree to 5-strongly agree were mostly positive, see Figure 5 for a breakdown of the results. Participants found the timeline the easiest view to understand (Mdn=4) and use (Mdn=4). Participants mostly agreed that the views helped accomplish the tasks (Mdn=5), that they would be useful within their writing process (Mdn=4), and that they would easily be integrated into their writing process (Mdn=4).
@@ 6. Discussion
We first discuss our results in light of participants’ comments before opening the discussion to other themes stemming from our observations and discussions with participants.
Visual writing can help find scenes and keep track of events, entities, and locations
Most participants agreed that “The visual representation was easy to understand” for all three views. Particularly, they mentioned how the views helped them. P5 appreciated that you can “click on a person, click on a location, and then see stuff pop up” instead of “having to look through […] many paragraphs […] or doing keyword search”. P4 explained “Just having a visual representation helps, you know, just with the different perspectives”. P3 added “Moving around the timeline was a good way of, kind of figuring out different events”. P1 and P5 also mentioned how they struggle to keep track of everything while writing their stories and that the system would help them keep track of that. For instance, P5 details that currently they “have booklets everywhere, and [they] have it annotated weird […] it’s a mess and this is so much easier. You got like, you know, your locations, your people […] keeping track of things. Organizing who’s where, what’s going on, being able to scroll through, like highlighting part of the text shown on the timeline […], making it easy to understand”.
Visual writing can help specify temporal, spatial, and entity-related edits
Participants reported high success rates when accomplishing the tasks and provided high ratings for statements such as “The visual representation helped me accomplish the tasks” and “The visual representation was easy to use”. This suggests that visual writing facilitates temporal, spatial, and entity-related edits through the views provided in our system. Participant’s comments further support these results. For example, P2 mentioned “If I add in an extra character, it gives you just enough text to build that story with that new character”. P4 appreciated that it made the story consistent when editing locations: “When I was dealing with the change where the hat was supposed to be left at home, the fact that it also updated the sentence where the hat flew off his head that is extremely useful because that’s the sort of inconsistencies that can easily creep in when you’re making modifications like that”. Similarly, P6 explained, “I really liked the ability to create relations between characters […] It’s something that made it a lot easier than any sort of other systems that I’ve used” and later added “[the tool] could be useful as a touch-up tool for when I want to make larger sweeping changes to a very long existing story […] I can confirm whether or not I want to keep each of these edits.”. This idea of reviewing the changes was important for participants, as P3 explained “You’d still have to review it, but it made enough changes that it [the story] was able to flow well.”
Visual writing can help explore and be creative
Many participants felt the tool was useful as a creative assistant. P6 explained “A lot of times, I am interested in how making a change to a given detail in a story would look […] I was appreciating it [the tool] as sort of an exploratory tool where even if it doesn’t end up producing something that I would personally publish, I think it’s really good to give an example that I can then engage with”. Similarly, P5 mentioned “You can just like, throw in a verb or the general action, interaction between them and it was kind of like spitballing ideas”. Regarding the timeline, P7 felt that “it’s an easy way to explore your options of what it sounds like if something happens here or if it sounds better when it’s there.”. They also mentioned how just looking at the visualization could help, for instance: “In terms of ideas, I could see how looking at the map, something could pop up into your head […] once you have, like, those objects laid out on a map, then that can help you spark inspiration”.
6.1. Observations and Participants’ Comments
6.1.1. Participants had different preferences
For instance, P4 mentioned “I really really like having a visual representation of the story, but this is not the visual representation I would choose […] I would like a combination of those two things [timeline and entity view] where those interlinks and connections appear within the timeline.”. Similarly, P6 explained how a single timeline would not be enough for them: “my brain interprets timelines in stories where I enjoy having things happening in multiple places at the same time and having a single linear approach makes it very hard to have that sort of interesting kind of development”. Other than visual representations, P4 also mentioned they “don’t want the system overwriting […] because then it’s not me expressing myself through text” and instead, they would prefer “just highlighting the places where things need to change, or the things are inconsistent”. This corroborates previous research on creative support tools needing to offer “many paths and many styles” to do tasks because users are unpredictable and have different preferences (Shneiderman et al., 2006).
6.1.2. Participants would have liked more control in using the visual representations to edit the story
While participants could express their intents with the tool, there are nuances in how an idea can be expressed in terms of writing style and words used, which is not currently supported by our system. For example, when discussing adding actions by connecting two entities, P8 mentioned “I feel like that’s very limiting how you wanted to express what you want versus just typing it out in a sentence because it kind of format it in a specific manner”. Similarly, P7 explained “in terms of expressing myself, I would say that it does help […] but I don’t think it allows for like total freedom of expression […] if you had something like, in your mind of how you wanted to say it, it might not come out that way”.
@@ 6.2. Limitations
6.2.1. The participants in our study were experienced and hesitant to using AI
@@
All the participants in our study reported having multiple years of creative writing experience. Similarly, on a scale from 1-never to 5-always, participants rated their frequency of use of AI services a median of one, with five participants reporting never using AI services. P4 explained “I write because I want to, so if I want to be writing then there’s ideas that I have, and I want to express them. I don’t want a system to tell me what it is that I am going to be expressing”. This confirms previous findings that writers have different opinions towards AI (Hertzmann, 2018; Biermann et al., 2022) and those opinions might have impacted our results.
6.2.2. The system and the representations do not scale to long stories
One main limitation of the representations used in the system is that they quickly become too crowded and unusable for long stories. Currently, the system solves this problem through filtering mechanisms when the cursor hovers over a sentence or a scene. This results in hiding the elements irrelevant to the passage currently being selected. But this only works for local edits. When the edits need to happen between scenes far apart in the story, then it becomes challenging to find the right scene or the correct location. We believe that, for these cases, new aggregated representations and semantic zooming would be necessary.
@@ 6.2.2. The system and the representations do not scale to long stories
6.2.3. The system might not have been robust enough to test the full potential of visual writing
@@
While the system was optimized to be robust, it still depends on a large language model to understand the story and the instructions properly. In our tests, but also during the study, we observed the system sometimes responding unexpectedly by modifying elements it should not or by refusing to do certain modifications. Additionally, despite our best efforts to parallelize requests and speed up the execution, some manipulations were still slow to execute (up to 10-15 seconds at times). These issues might have confounded the results of our study. For example, it is known that to promote user exploration, interactions should be fast and with effects immediately visible (Masson et al., 2022; Shneiderman et al., 2006; Shneiderman, 1983).
6.3. Future Work
6.3.1. Visual writing to edit writing style and story plots
Creative writing is more than events, actors, locations, and time. We believe that visual writing could also help editing other aspects, such as the style of the writing. As proposed by Kempen et al. (1987), the syntax tree of the sentences could be made visible and be manipulable to restructure sentences. Alternatively, the tone, the length of the sentences, or the sentence structure could be represented using histograms or by embellishing the text. Then, manipulating these visuals would allow tweaking the style of the story. At a higher level, visual writing could also help revise the plot of the story. Similar to work on manipulating outlines (Dang et al., 2022a; Zhang et al., 2023), the idea would be to represent the main plot events of the story and then make them manipulable to reorder or change them.
6.3.2. Exploring new visual interactions
The idea of visual writing is broader than the representations and interactions proposed by our system. Any visual representation better suited for a task could be beneficial. For example, one could imagine using images to give the system an indication of what an object, character, or location look like. Then, the image could be modified to update the story (e.g., painting the hair of a character blond would edit the story so that the description matches the new hair colour). The system could also allow highlighting part of the image to indicate which aspect should be more or less emphasized in the description. Similarly, while our implementation uses LLM, visual writing could be implemented without and made more or less autonomous (Morris et al., 2024). Using LLM comes about with many challenges related to biases, privacy, and plagiarism (Weidinger et al., 2021; Meyer et al., 2023). An alternative, suggested by P4, could be to have the visual writing system only highlight the parts that should be modified to help the writer identify and edit them themselves.
6.3.3. Suggestions and creativity support in the visual space
The mental processes involved in interpreting visuals differ from those involved in processing language (Paivio, 1990). Future work could explore the impact on creativity of writing stories visually rather than with text. Similarly, traditional creativity support techniques for writers would be interesting to adapt to the visual space. For instance, there are many benefits of phrase suggestions (Buschek et al., 2021; Bhat et al., 2023), and future work could investigate if there are similar benefits to suggestions in the visual space. These visual suggestions would be shown when an interaction with a visual element is initiated to highlight different alternatives and to “auto-complete” the interaction.
@@ 7. Conclusion
We proposed visual writing, an approach to writing by manipulating visual representations to express editing intents. Through a taxonomy of possible visual representations based on narratology and a prototype implementation of some of these visual representations, we demonstrated one possible design for visual writing tools. A study with eight creative writers showed the potential of creative writing to help keep track of story elements, rapidly specify edits, and explore story variations in a way that encourages creativity. Broadly, our work advocates for a new generation of writing support tools that allow text editing through direct manipulation of visual representations.
@@ Acknowledgements.
This work was supported by NSERC under grant RGPIN-2018-05072.References
@@
- Arnold et al. (2021) Kenneth C Arnold, April M Volzer, and Noah G Madrid. 2021. Generative Models Can Help Writers without Writing for Them. In IUI Workshops.
@@
- Bhat et al. (2023) Advait Bhat, Saaket Agashe, Parth Oberoi, Niharika Mohile, Ravi Jangir, and Anirudha Joshi. 2023. Interacting with Next-Phrase Suggestions: How Suggestion Systems Aid and Influence the Cognitive Processes of Writing. In Proceedings of the 28th International Conference on Intelligent User Interfaces (IUI ’23). Association for Computing Machinery, New York, NY, USA, 436–452. https://doi.org/10.1145/3581641.3584060
- Biermann et al. (2022) Oloff C. Biermann, Ning F. Ma, and Dongwook Yoon. 2022. From Tool to Companion: Storywriters Want AI Writers to Respect Their Personal Values and Writing Strategies. In Designing Interactive Systems Conference. ACM, Virtual Event Australia, 1209–1227. https://doi.org/10.1145/3532106.3533506
@@
- Burnett and McIntyre (1995) Margaret M. Burnett and David W. McIntyre. 1995. Visual Programming. Computer 28, 03 (March 1995), 14–16. https://doi.org/10.1109/MC.1995.10027
- Buschek et al. (2021) Daniel Buschek, Martin Zürn, and Malin Eiband. 2021. The Impact of Multiple Parallel Phrase Suggestions on Email Input and Composition Behaviour of Native and Non-Native English Writers. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3411764.3445372
- Calderwood et al. (2020) Alex Calderwood, Vivian Qiu, Katy Ilonka Gero, and Lydia B. Chilton. 2020. How Novelists Use Generative Language Models: An Exploratory User Study.. In HAI-GEN+ User2agent@ IUI.
- Carroll et al. (2009) Erin A. Carroll, Celine Latulipe, Richard Fung, and Michael Terry. 2009. Creativity factor evaluation: towards a standardized survey metric for creativity support. In Proceedings of the Seventh ACM Conference on Creativity and Cognition (Berkeley, California, USA) (C&C ’09). Association for Computing Machinery, New York, NY, USA, 127–136. https://doi.org/10.1145/1640233.1640255
- Chakrabarty et al. (2024) Tuhin Chakrabarty, Vishakh Padmakumar, Faeze Brahman, and Smaranda Muresan. 2024. Creativity Support in the Age of Large Language Models: An Empirical Study Involving Emerging Writers. arXiv:2309.12570 [cs]
- Cherry and Latulipe (2014) Erin Cherry and Celine Latulipe. 2014. Quantifying the Creativity Support of Digital Tools through the Creativity Support Index. ACM Trans. Comput.-Hum. Interact. 21, 4 (June 2014), 21:1–21:25. https://doi.org/10.1145/2617588
- Chung et al. (2022) John Joon Young Chung, Wooseok Kim, Kang Min Yoo, Hwaran Lee, Eytan Adar, and Minsuk Chang. 2022. TaleBrush: Sketching Stories with Generative Pretrained Language Models. In CHI Conference on Human Factors in Computing Systems. ACM, New Orleans LA USA, 1–19. https://doi.org/10.1145/3491102.3501819

### v2

@@ 6.2. Results
Participants completed the tasks and the free-form exploration without major difficulties. Below, we summarize the quantitative findings using descriptive statistics. Following recommendations for similar study designs (Zhu and Kolassa, 2018; Masson et al., 2023), confidence intervals use the studentized bootstrap with 10,000 repetitions. Themes are based on their relevance to our research questions, as per an interpretivist approach to reflexive thematic analysis (Braun and Clarke, 2022).
@@
Finding scenes and keeping track of events, entities, and locations. As a search tool, the results of this study corroborated those of the first study. Most participants agreed that “The visual representation was easy to understand” for all three views (Figure 7). W5 appreciated that you can “click on a person, click on a location, and then see stuff pop up” instead of “having to look through […] many paragraphs […] or doing keyword search”. W4 explained “Just having a visual representation helps, you know, just with the different perspectives”. W3 added “Moving around the timeline was a good way of, kind of figuring out different events”. W1 and W5 also mentioned how they usually struggle to keep track of everything while writing their stories and that the system could help them. For instance, W5 details that currently they “have booklets everywhere, and [they] have it annotated weird […] it’s a mess and this is so much easier. You got like, you know, your locations, your people […] keeping track of things. Organizing who’s where, what’s going on, being able to scroll through, like highlighting part of the text shown on the timeline […] making it easy to understand”. This was also reflected in participants reporting low mental demand on the NASA-TLX (M=7.49 SD=1.95).
Specifying temporal, spatial, and entity-related edits. On a 5-point semantic differential scale from 1-unsuccessful to 5-successful, participants rated their success M=4.25 (95% CI: [3.8, 4.6]) at accomplishing the tasks. They were most successful at the task about adding and removing entities (M=4.5, 95% CI: [3.8, 5]), changing the locations of entities (M=4.25, 95% CI: [3.4, 4.8]), and reordering events (M=4, 95% CI: [2.8, 4.7]). Participants used an average of 2.42 visual edits to accomplish a task. In the free-form part, the most common visual edits were adding actions (36%), editing actions (26%), and editing traits (19%). Specifically, participants appreciated that the edits produced text as a good starting point. W2 mentioned “If I add in an extra character, it gives you just enough text to build that story with that new character”. W4 appreciated that it fixed inconsistencies when editing locations: “the fact that it also updated the sentence where the hat flew off his head that is extremely useful because that’s the sort of inconsistencies that can easily creep in when you’re making modifications like that”. Similarly, W6 expressed, “I really liked the ability to create relations between characters”.
On the NASA-TLX, they also reported high-performance (M=7.49 SD=1.95) for low effort (M=3.39, SD=2.17), frustration (M=2.5, SD=1.2), temporal demand (M=1.18, SD=1.04), and physical demand (M=1.18, SD=1.04). For example, W6 further expressed that “[the tool] could be useful as a touch-up tool for when I want to make larger sweeping changes to a very long existing story […] I can confirm whether or not I want to keep each of these edits” stating that the system “made it a lot easier than any sort of other systems that I’ve used”. The idea of reviewing the changes was important for participants, as W3 explained “You’d still have to review it, but it made enough changes that it [the story] was able to flow well.”
Helping explore and be creative. Participants generated an average CSI score of 71.5 (SD=21.12) (Table 2). They felt the tool was useful as a creative assistant, even when the text produced was not up to their standards. W6 explained “A lot of times, I am interested in how making a change to a given detail in a story would look […] I was appreciating [the tool] as sort of an exploratory tool where even if it doesn’t end up producing something that I would personally publish, I think it’s really good to give an example that I can then engage with”. Similarly, W5 mentioned “You can just like, throw in a verb or the general action, interaction between them and it was kind of spitballing ideas”. Regarding the timeline, W7 felt that “it’s an easy way to explore your options of what it sounds like if something happens here or if it sounds better when it’s there.”. They also mentioned how just looking at the visualization could help, for instance: “In terms of ideas, I could see how looking at the map, something could pop up into your head […] once you have, like, those objects laid out on a map, then that can help you spark inspiration”.
Impact on the writing workflow. Perhaps due to the study setting, all participants heavily used the system during the free-form part. Most participants used the system to explore, “just to see what would happen” (W2), then read the suggested changes and kept them or undid them either by editing an event (W2, W6), undoing using the history graph (W3), deleting actions and characters (W1, W4, W6), or rewriting passages manually (W4, W8). When exploring, participants mostly used the actions and the location views. The timeline was only used to select specific scenes to edit. Participants also had different strategies when creating events, for example, W1, W4, and W7 wrote longer and detailed actions such as “Interrupts conversations, she says that she saw it too, and she is also confused” (W1) or explicitly writing a dialogue “Says ‘what are you doing here?”’ (W7), whereas others wrote one or two-word actions or descriptions. W8 had the most unique workflow in that they used the visuals to edit the text only once. Instead, they spent most of the session writing their story manually and would periodically refresh the visualization based on what they wrote and then reflect. They explained “In my most focused state I would just be focused on writing, but then when I was done, I was out of it, I would scan back and forth between text and the [visuals]”.
Mismatches between the system and participants’ existing workflows. Questions related to integrating the system in the writing process were rated the lowest (but still with an Mdn=4). Similar to the findings from the first study, the most critical participants explained that the visuals did not match their workflow and the visualizations they already used. W4 mentioned “I really really like having a visual representation of the story, but this is not the visual representation I would choose […] I would like a combination of those two things [timeline and entity view] where those interlinks and connections appear within the timeline.”444While this visualization was not implemented in our system, it can be modelled with our framework: start with characters, connect by events, position by locations, and unfold by time. Similarly, W6 explained how a single timeline would not be enough for them: “my brain interprets timelines in stories where I enjoy having things happening in multiple places at the same time and having a single linear approach makes it very hard to have that sort of interesting kind of development”555Again, this can be modelled by our framework by unfolding by both time and temporality to visualize nonlinear narratives. Other than visual representations, W4 also mentioned they “don’t want the system overwriting […] because then it’s not me expressing myself through text” and instead, they would prefer “just highlighting the places where things need to change, or the things are inconsistent”.
Participants would have liked more control in using the visual representations to edit the story. While participants could express their intents with the tool, there are nuances in how an idea can be expressed in terms of writing style and voice, which is not currently supported by our system. For example, when discussing adding actions by connecting two entities, W8 mentioned “I feel like that’s very limiting how you wanted to express what you want versus just typing it out in a sentence because it kind of format it in a specific manner”. Similarly, W7 explained “in terms of expressing myself, I would say that it does help […] but I don’t think it allows for like total freedom of expression […] if you had something, in your mind of how you wanted to say it, it might not come out that way”.
@@ 7. Discussion
We propose visual story-writing, an approach that uses visual representations of story elements to support writing and revising narrative texts. Despite how little writing support tools have used visualizations (Zhao et al., 2025), our findings show that they hold great promise to help review and ideate stories. The space of possibilities is large, and our framework of story constructs is an initial attempt at structuring it. In our study, we looked at a subset of this space, using simple representations that did not cover all story elements. However, we encourage others to continue this line of research.
One question with tools like this is how it might change the role of writers. AI-assisted writing tools tend to make us write less and review and edit more (Buschek, 2024). We believe visual story-writing can accommodate both traditional and “editorial” writing roles: in its simplest form, visual story-writing is automatizing the creation of visualizations writers were already using (Zhao et al., 2025; Ackerman, 2017; Neuwirth and Kaufer, 1989; SaveTheCat, 2014; Articy, 2014), with the added bonus of serving as input, if the writer desires it.
Another question is how visual story-writing could be integrated with other writing support tools. We view visual story-writing as complementary to tools that rely on textual prompts (e.g., ChatGPT or Grammarly). Like work on direct manipulation with LLMs has shown (Masson et al., 2024), we expect visual story-writing to be preferable when writers have a clearer idea of the edit they want to perform (i.e., “move this character there”) because it helps refer to story elements unambiguously. Whereas textual prompts are preferable when the goal is vague (e.g., “make this story sound better”).
Finally, our work targets the creative writing of narratives that involve tangible elements like characters and locations. It applies not only to novels but also to screenwriting for films, TV shows, video games, and theatre plays. However, it is unclear how visual story-writing could work in domains like academic writing, where the elements manipulated are not so tangible.
7.1. Future Work
7.1.1. Extending the framework and exploring new visual interactions
Our framework (§3) operates at a different conceptual level than typical visualization frameworks, such as the grammar of graphics (Wilkinson et al., 2005). It combines story elements into constructs that are then linked to visuals without prescribing how to extract story elements or what the visualizations should look like, allowing more freedom in implementation. For example, our system represents locations as bubbles instead of a traditional map. Alternatively, images could indicate what an object, character, or location looks like, allowing manipulation (e.g., changing a character’s hair colour) or emphasis on certain descriptive points. More work is needed to identify which constructs can best serve writers within visual story-writing workflows, and how to best represent them. For example, inter-character relationships are important in a story (Bochner et al., 1997) and a view constructed by connecting actors would help visualize and edit these relationships. Additionally, the framework can be extended to include story elements beyond those highlighted by Bal (Bal, 1997). For example, ”emotions” could be a basic building block, as used in StoryPrint (Watson et al., 2019), and ”motivations” could be another, as suggested by P11 during the first study. In fact, other narratology models, such as Propp’s morphology of the folk tale (Propp and Pirkova-Jakobson, 2010), Todorov’s principles of narratives (Todorov, 1971), or the contemporary concepts on narrative time by Hume (Hume, 2014) would lead to a different framework.
7.1.2. Visual editing of writing style and story plots
Creative writing encompasses more than just events, actors, locations, and time. The approach could also help edit other aspects, such as writing style. As proposed by Kempen et al. (1987), manipulating the syntax tree can restructure sentences. Alternatively, tone, sentence length, or structure could be represented with histograms or text embellishments, allowing style adjustments through visual manipulation. At a higher level, visual story-writing can aid revising the plot. Similar to work on manipulating outlines (Dang et al., 2022; Zhang et al., 2023), main plot events can be visually represented to be reordered or changed.
7.1.3. Suggestions and creativity support in the visual space
The mental processes involved in interpreting visuals differ from those involved in processing language (Paivio, 1990). Future work could explore the impact on the creativity of writing stories through visuals rather than with text. Similarly, typical writing support features would be interesting to adapt to the visual space. For instance, there are many benefits of phrase suggestions (Buschek et al., 2021; Bhat et al., 2023), and future work could investigate if there are similar benefits to suggestions in the visual space. These visual suggestions would essentially try to “auto-complete” an interaction initiated on a visual element.
7.1.4. Supporting long stories
For long stories, the visuals have to find the right balance between overview and details. Our prototype attempts to show the details of all the events and entities and rely on interactions to declutter the views, by merging overlapping actions and allowing users to hover or select a passage to filter what is shown. If details are prioritized, this view could be unfolded by temporality to “spread out” the interactions over time. Unfolding again by time would visualize complex nonlinear narratives. Alternatively, interaction techniques such as details-on-hover or semantic zooming could also help manage longer stories.
7.1.5. Visualization-builder to support custom visuals
The studies showed creative writers have various preferences in terms of visuals. One way to support these workflows would be to help users construct their own visuals. Our procedural framework could inform such a view-builder: the eight story elements would be the basic building blocks, and users could construct a view by applying the different operators. Once the view is constructed, it could become interactive and automatically populated from the text.
@@ 7.2. Limitations
7.2.1. Participants’ experience and reservation towards AI might have impacted our results
@@
In study 1, half the participants had no creative writing experience, meaning they were possibly at a disadvantage to accomplish the tasks. Conversely, all participants in study 2 had years of creative experience, and, while they could comment on how visual story-writing could be integrated in their workflows, the results could not tell us about the benefits of the tool for beginners. Similarly, some participants expressed reservations towards AI and reported rarely, if ever, using it. W4 explained “I write because I want to, so if I want to be writing then there’s ideas that I have, and I want to express them. I don’t want a system to tell me what it is that I am going to be expressing”. Considering writers’ have strong opinions about AI (Hertzmann, 2018; Biermann et al., 2022), this might have biased the results.
@@
7.2.2. The system might not have been robust enough to test the full potential of visual story-writing
@@
The system relied on an LLM that occasionally responded unexpectedly, such as modifying unintended elements or refusing to do certain modifications. Additionally, some manipulations were slow to execute (up to 10-15 seconds) despite our efforts to parallelize requests. These issues may have affected our results as interactions should be seamless and paired with immediate effects to promote user exploration (Masson et al., 2022; Shneiderman et al., 2006; Shneiderman, 1983).
@@ 8. Conclusion
We proposed visual story-writing, an approach to support writing by reviewing and manipulating visual representations. In doing so, we defined a framework to help inform the design of story visualizations. By applying this framework, we implemented a prototype system demonstrating one possible design for visual story-writing tools. Two studies covering the different aspects of the writing process showed the potential of creative writing to help keep track of story elements, rapidly specify edits, and explore story variations in a way that encourages creativity. Broadly, our work advocates for a new generation of writing support tools that embed visualization to help review and edit text narratives.
@@ Acknowledgements.
We thank Daniel Aureliano Newman for his help revising the questions from study 1. This project was undertaken thanks to funding from IVADO, the Canada First Research Excellence Fund, and NSERC Discovery Grant RGPIN-2018-05072.References
@@
- Arnold et al. (2021) Kenneth C Arnold, April M Volzer, and Noah G Madrid. 2021. Generative Models Can Help Writers without Writing for Them. In Joint Proceedings of the ACM IUI 2021 Workshops. CEUR, College Station, USA, 8.
- Articy (2014) Articy Software GmbH & Co 2014. Articy: narrative design for interactive projects. Articy Software GmbH & Co. https://www.articy.com
@@
- Bal (2021) Mieke Bal. 2021. Narratology in Practice. University of Toronto Press, Toronto.
- Bensaid et al. (2021) Eden Bensaid, Mauro Martino, Benjamin Hoover, and Hendrik Strobelt. 2021. FairyTailor: A Multimodal Generative Framework for Storytelling. arXiv:2108.04324 [cs] doi:10.48550/arXiv.2108.04324
- Bhat et al. (2023) Advait Bhat, Saaket Agashe, Parth Oberoi, Niharika Mohile, Ravi Jangir, and Anirudha Joshi. 2023. Interacting with Next-Phrase Suggestions: How Suggestion Systems Aid and Influence the Cognitive Processes of Writing. In Proceedings of the 28th International Conference on Intelligent User Interfaces (IUI ’23). Association for Computing Machinery, New York, NY, USA, 436–452. doi:10.1145/3581641.3584060
- Biermann et al. (2022) Oloff C. Biermann, Ning F. Ma, and Dongwook Yoon. 2022. From Tool to Companion: Storywriters Want AI Writers to Respect Their Personal Values and Writing Strategies. In Designing Interactive Systems Conference. ACM, Virtual Event Australia, 1209–1227. doi:10.1145/3532106.3533506
- Bochner et al. (1997) Arthur P. Bochner, Carolyn Ellis, and Lisa M. Tillman-Healy. 1997. Relationships as Stories. In Handbook of Personal Relationships: Theory, Research and Interventions, 2nd Ed. John Wiley & Sons, Inc., Hoboken, NJ, US, 307–324.
@@
- Braun and Clarke (2022) Virginia Braun and Victoria Clarke. 2022. Thematic Analysis: A Practical Guide. SAGE Publications Ltd, Los Angeles.
- Bressa et al. (2024) Nathalie Bressa, Jordan Louis, Wesley Willett, and Samuel Huron. 2024. Input Visualization: Collecting and Modifying Data with Visual Representations. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 499, 18 pages. doi:10.1145/3613904.3642808
- Burnett and McIntyre (1995) Margaret M. Burnett and David W. McIntyre. 1995. Visual Programming. Computer 28, 03 (March 1995), 14–16. doi:10.1109/MC.1995.10027
- Buschek (2024) Daniel Buschek. 2024. Collage Is the New Writing: Exploring the Fragmentation of Text and User Interfaces in AI Tools. In Proceedings of the 2024 ACM Designing Interactive Systems Conference (DIS ’24). Association for Computing Machinery, New York, NY, USA, 2719–2737. doi:10.1145/3643834.3660681
- Buschek et al. (2021) Daniel Buschek, Martin Zürn, and Malin Eiband. 2021. The Impact of Multiple Parallel Phrase Suggestions on Email Input and Composition Behaviour of Native and Non-Native English Writers. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–13. doi:10.1145/3411764.3445372
- Calderwood et al. (2020) Alex Calderwood, Vivian Qiu, Katy Ilonka Gero, and Lydia B. Chilton. 2020. How Novelists Use Generative Language Models: An Exploratory User Study.. In IUI ’20 Workshops. CEUR, Cagliari, Italy, 5.
- Carroll et al. (2009) Erin A. Carroll, Celine Latulipe, Richard Fung, and Michael Terry. 2009. Creativity factor evaluation: towards a standardized survey metric for creativity support. In Proceedings of the Seventh ACM Conference on Creativity and Cognition (Berkeley, California, USA) (C&C ’09). Association for Computing Machinery, New York, NY, USA, 127–136. doi:10.1145/1640233.1640255
- Cherry and Latulipe (2014) Erin Cherry and Celine Latulipe. 2014. Quantifying the Creativity Support of Digital Tools through the Creativity Support Index. ACM Trans. Comput.-Hum. Interact. 21, 4 (June 2014), 21:1–21:25. doi:10.1145/2617588
- Chung et al. (2022) John Joon Young Chung, Wooseok Kim, Kang Min Yoo, Hwaran Lee, Eytan Adar, and Minsuk Chang. 2022. TaleBrush: Sketching Stories with Generative Pretrained Language Models. In CHI Conference on Human Factors in Computing Systems. ACM, New Orleans LA USA, 1–19. doi:10.1145/3491102.3501819
- Chung et al. (2025) John Joon Young Chung, Melissa Roemmele, and Max Kreminski. 2025. Toyteller: AI-powered Visual Storytelling Through Toy-Playing with Character Symbols. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, 1–23. doi:10.1145/3706598.3713435
- Cinnamon (2017) Nadieh Bremer-Visual Cinnamon. 2017. All Fights in Dragon Ball Z. https://dragonballz.visualcinnamon.com.